# Denoising weak lensing mass maps with diffusion model: systematic comparison with generative adversarial network

### v1

@@ A.1 Networks in GAN
For GAN, the generator employs U-Net architecture (Ronneberger et al. 2015), and the discriminator utilises four convolutional blocks. U-Net consists of eight convolutional blocks, starting with 64 channels and increasing to 512 channels in the deepest layer. Each convolutional block performs convolution (or transposed convolution) with a kernel size of $4\times 4$. The activation function used in the encoder part is LeakyReLU, in the decoder part it is ReLU, and tangent hyperbolic function is applied in the final layer. Additionally, dropout with a rate of $0.5$ is applied to intermediate layers. In the discriminator network, all convolutional layers use a $4\times 4$ kernel, and the activation function is LeakyReLU. The initial number of channels is 2, combining the input image and the target image. As the layers progress, the number of channels increases up to 512, and finally, it reduces to 1 channel. The discriminator takes only a part of the entire image to capture only high-frequency modes, which is referred to as PatchGAN in Isola et al. (2017). We employ $30\times 30$ PatchGAN in the discriminator. The batch normalisation is applied in both networks. The total number of parameters of networks is $5.4\times 10^{7}$ in the generator and $2.8\times 10^{6}$ in the discriminator.
@@ A.2 Networks in DM
DM also adopts U-Net architecture, and this U-Net consists of three downsampling and upsampling layers. The input has 64 channels, and the number of channels doubles at each downsampling layer, reaching 512 at the intermediate layer. Each layer consists of two Residual Network (ResNet; He et al. 2015) blocks with a kernel size of $3\times 3$, a self-attention layer, and a downsampling or upsampling operation. The ResNet blocks employ group normalisation, SiLU activation function, and residual convolution. The bottleneck processes the input sequentially through a ResNet block, a self-attention layer, and ResNet block. The total number of parameters of the network is $6.3\times 10^{7}$.
Appendix B Direct and indirect denoising
@@
Here, we present the results of the comparison between the target data in the denoising process. We define direct denoising as the case when the target field of GAN and DM is the noiseless convergence map. On the other hand, in indirect denoising, the models learn the mapping from the noisy map to the noise map itself. Then, the noiseless mass map can be estimated by subtracting the predicted noise map from the noisy map. Among statistics, the power spectra and PDFs are shown in Figures 16 and 17, respectively For indirect denoising, we use the same hyper-parameters as the direct denoising. In general, it turns out that direct denoising performs better than indirect denoising. We conclude that directly estimating the noiseless convergence map yields better results. However, Shirasaki et al. (2019) have reported cases where the indirect transformation achieves higher accuracy, suggesting that the optimal approach may depend on the data set.
@@
Acknowledgements.
This work was supported in part by JSPS KAKENHI Grant Number JP24H00215, JP25H01513, JP25K17380 (K.O.) and JP24H00221 (M.S.). This research was conducted using the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, The University of Tokyo.

### v2

@@ A.1 Networks in GAN
For GAN, the generator employs U-Net architecture (Ronneberger2015), and the discriminator utilises four convolutional blocks. U-Net consists of eight convolutional blocks, starting with 64 channels and increasing to 512 channels in the deepest layer. Each convolutional block performs convolution (or transposed convolution) with a kernel size of $4\times 4$. The activation function used in the encoder part is LeakyReLU, in the decoder part it is ReLU, and the tangent hyperbolic function is applied in the final layer. Additionally, dropout with a rate of $0.5$ is applied to intermediate layers. In the discriminator network, all convolutional layers use a $4\times 4$ kernel, and the activation function is LeakyReLU. The initial number of channels is 2, combining the input image and the target image. As the layers progress, the number of channels increases up to 512, and finally, it reduces to 1 channel. The discriminator takes only a part of the entire image to capture only high-frequency modes, which is referred to as PatchGAN in Isola2017. We employ $30\times 30$ PatchGAN in the discriminator. The batch normalisation is applied in both networks. The total number of parameters of networks is $5.4\times 10^{7}$ in the generator and $2.8\times 10^{6}$ in the discriminator.
@@ A.2 Networks in DM
DM also adopts U-Net architecture for the reverse process, and this U-Net consists of three downsampling and upsampling layers. The input has 64 channels, and the number of channels doubles at each downsampling layer, reaching 512 at the intermediate layer. Each layer consists of two Residual Network (ResNet; He2015) blocks with a kernel size of $3\times 3$, a self-attention layer, and a downsampling or upsampling operation. The ResNet blocks employ group normalisation, SiLU activation function, and residual convolution. The bottleneck processes the input sequentially through a ResNet block, a self-attention layer, and ResNet block. The total number of parameters of the network is $6.3\times 10^{7}$.
Appendix B Denoising performance in the high source number density case
So far, we have assumed the source number density which mimics Subaru HSC WL survey. We demonstrate our denoising models are effective with a higher source number density, i.e., the case with lower shape noise amplitude. To this end, we repeat our denoising process with higher source number density. Here, the source number density is adopted as $n_{\mathrm{g}}=64.2\,\mathrm{arcmin}^{-2}$, which is determined based on COSMOS catalogues (Schrabback2010). We train and test both GAN and DM in the same configuration as our fiducial mock maps, i.e., $39,000$ maps for training and $1,000$ maps for testing, but GAN is trained only once instead of five in the fiducial model. The results of RMSE and Pearson correlation coefficient $\rho$ are summarised in Table B. In comparison with the fiducial result shown in Table 4.3, the similar level of denoising performance is confirmed with the higher source number density.
[…]
Appendix C Noise schedule of diffusion model
Here, we address how the noise schedule $\alpha_{t}(=1-\beta_{t})$ (Eq. 16) affects the reconstructed cosmological statistics. Three different noise schedules are considered: linear schedule, quadratic schedule, and cosine schedule. The linear and quadratic schedules, which are denoted as $\beta^{\mathrm{linear}}_{t}$ and $\beta^{\mathrm{quad}}_{t}$, respectively, start from $\beta_{1}=10^{-6}$ to $\beta_{T}=0.01$ are defined as
$\beta^{\mathrm{linear}}_{t}$ $=\beta_{1}(1-t)+\beta_{T}t,$ (44)
$\beta^{\mathrm{quad}}_{t}$ $=\left(\sqrt{\beta_{1}}(1-t)+\sqrt{\beta_{T}}t\right)^{2}.$ (45)
The cosine schedule, which is originally proposed by Nichol2021, is designed to prevent the diffusion process from being noise-dominated within the first few steps, which allows for more efficient training. The cosine schedule is expressed as
$\beta^{\mathrm{cos}}_{t}$ $=\min\left(1-\frac{\bar{\alpha}^{\mathrm{cos}}_{t}}{\bar{\alpha}^{\mathrm{cos}}_{t-1}},0.999\right),$ (46)
$\bar{\alpha}^{\mathrm{cos}}_{t}$ $=\frac{f(t)}{f(0)},$ (47)
$f(t)$ $=\cos\left(\frac{t/T+s}{1+s}\cdot\frac{\pi}{2}\right)^{2},$ (48)
where $s=8\times 10^{-3}$ is a parameter introduced to prevent $\beta_{t}$ from becoming too small.
The reconstructed power spectra, PDFs, peak and minima counts with three noise schedules are shown in Figures 16 and 17, Except PDFs, where the linear and cosine schedules yield slightly better reconstruction performance than the quadratic schedule, the quadratic noise schedule reproduces the other statistics more accurately, especially at small scales and in low-$\mathrm{SNR}$ regions. This feature can be explained as the quadratic schedule takes more steps with smaller noise amplitudes and high-frequency modes can be more efficiently learned.
Figure 16: The comparison of the power spectra and the PDF of denoised maps with DM between the three noise schedules: quadratic (orange), linear (green) and cosine (red). The solid lines of the linear and cosine schedules are obtained by a single sampling, while the quadratic schedule curves are the same as those in Figure 5 and 10
Figure 17: Same as Figure 16 but for peak and minima counts.
Appendix D Direct and indirect denoising
@@
Here, we present the results of the comparison between the target data in the denoising process. We define direct denoising as the case when the target field of GAN and DM is the noiseless convergence map. On the other hand, in indirect denoising, the models learn the mapping from the noisy map to the noise map itself. Then, the noiseless mass map can be estimated by subtracting the predicted noise map from the noisy map. Among statistics, the power spectra and PDFs are shown in Figures 18 and 19, respectively For indirect denoising, we use the same hyper-parameters as the direct denoising. In general, it turns out that direct denoising performs better than indirect denoising. We conclude that directly estimating the noiseless convergence map yields better results. However, Shirasaki2019 have reported cases where the indirect transformation achieves higher accuracy, suggesting that the optimal approach may depend on the data set.